OpenStudy (anonymous):

Solve for x. 3 – 2x(2 – 5) = 3(4 – x)

OpenStudy (anonymous):

First distribute -2x(2-5) and 3(4-x) using distributive property. 3-2x(2-5)=3(4-x) 3-2x*2-2x*(-5)=3*4-3*4x 3-4x+10x=12-12x

OpenStudy (anonymous):

Bring all x terms to one side and bring all the numbers to other side. Then solve for x.

OpenStudy (anonymous):

x then should = 1